### v1

# MessIRve: A Large-Scale Spanish Information Retrieval Dataset

Abstract
Information retrieval (IR) is the task of finding relevant documents in response to a user query. Although Spanish is the second most spoken native language, current IR benchmarks lack Spanish data, hindering the development of information access tools for Spanish speakers. We introduce MessIRve, a large-scale Spanish IR dataset with around 730 thousand queries from Google’s autocomplete API and relevant documents sourced from Wikipedia. MessIRve’s queries reflect diverse Spanish-speaking regions, unlike other datasets that are translated from English or do not consider dialectal variations. The large size of the dataset allows it to cover a wide variety of topics, unlike smaller datasets. We provide a comprehensive description of the dataset, comparisons with existing datasets, and baseline evaluations of prominent IR models. Our contributions aim to advance Spanish IR research and improve information access for Spanish speakers.
MessIRve: A Large-Scale Spanish Information Retrieval Dataset
Francisco Valentini1, Viviana Cotik1,2, Damián Furman1, Ivan Bercovich3, Edgar Altszyler4, Juan Manuel Pérez1 1CONICET-UBA. Instituto de Ciencias de la Computación. Buenos Aires, Argentina 2UBA. FCEyN. Departamento de Computación. Buenos Aires, Argentina 3University of California, Santa Barbara. California, U.S.A. 4Quantit {fvalentini,jmperez,vcotik}@dc.uba.ar, damian.a.furman@gmail.com, ibercovich@ucsb.edu, edgar@bequantit.com
1 Introduction
Given a user query, Information retrieval (IR) is the process of locating relevant documents from a collection of data. IR systems have become increasingly important as they are the backbone for the currently widespread Retrieval Augmented Generation (RAG) systems, in which IR provides relevant passages that are subsequently fed into a Large Language Model (LLM). This approach has been shown to improve the quality of generated text, reducing the likelihood of factual errors or hallucinations by LLMs in several tasks (Cheng et al., 2023; Jiang et al., 2023; Cheng et al., 2024; Lin et al., 2024).
Natural Language Processing (NLP) research relies heavily on evaluation datasets, which establish a common ground for comparing retrieval algorithms and guide the development of new methods.
While a wealth of resources is available for English, other languages often rely on multilingual models and datasets. When addressing tasks in Spanish, the resources are notably scarce. Spanish is the the second most spoken language in the world by number of native speakers, with over 486 million speakers. It is also the official language of 20 countries and is widely spoken in the United States (Eberhard et al., 2024).
Despite its prevalence, at the time of writing this paper, the retrieval section of the popular MTEB benchmark (Muennighoff et al., 2023) includes datasets for English, Chinese, French, and Polish, but omits Spanish.111huggingface.co/spaces/mteb/leaderboard
The lack of good quality evaluation datasets is a major challenge for the development of IR systems for Spanish. Which systems are best for Spanish IR? This key question remains unanswered due to the lack of a comprehensive Spanish IR benchmark.
In this work, we address this gap by:
  1. 1.
    Identifying and describing the Spanish IR datasets that are currently available.
  2. 2.
    Introducing MessIRve, a new large-scale dataset in Spanish that accounts for the varieties of Spanish spoken accross the Spanish-speaking countries. [^2] MessIRve means works for me in Spanish (“me sirve”). The reference to Lionel Messi, player of the most popular sport in Spanish-speaking countries, football, also stresses the importance of using topics that are relevant to Spanish speakers.
  3. 3.
    Providing baseline evaluations of relevant existing models on the new dataset.
  4. 4.
    Making the dataset publicly available so that other researchers can use it to develop and evaluate IR systems for Spanish. [^3] The dataset is available at https://huggingface.co/datasets/spanish-ir/google_qrels under an appropriate license.
These contributions aim to spur further research in Information Retrieval for the Spanish language and to facilitate the development of efficient information access tools for Spanish speakers.
2 Related work
Dataset #Docs Train Dev Test
#Q #J/#Q #Q #J/#Q #Q #J/#Q
mMARCO (Bonifacio et al., 2021) 8 841 823 532 761 1.00 - - 6 979 1.07
SQAC (Gutiérrez-Fandiño et al., 2022) 6 247 14 934 1.01 1 861 1.00 1 908 1.00
MIRACL (Zhang et al., 2023) 10 373 953 2 162 9.96 648 9.94 1 515 9.95
PRES (Kamateri et al., 2019) 10 037 - - - - 167 7.72
Multi-EuP (Yang et al., 2023) 2 371 - - - - 633 3.75
MessIRve (full) 14 047 759 571 120 1.05 - - 160 099 1.02
Table 1: Spanish IR Datasets. #: number of, Docs: documents, Q: queries, J: annotated relevance judgments. Judgments indicate positive (relevant) documents in all datasets, except for MIRACL, which includes negative (non-relevant) annotations as well. The full split of MessIRve includes queries that are not specific to any country plus the set of unique queries from all countries (see section 3.2).
In this section, we review the Spanish IR datasets that are well-documented and publicly accessible. Relevant statistics of these datasets are presented in Table 1.
Following standard terminology, we refer to a dataset as a collection of queries, documents, and relevance judgments (Thakur et al., 2021). A query is a user’s need for information, a document is a passage that may contain the information the user seeks, and a relevance judgment indicates whether a document is relevant to a query. The collection of all documents is referred to as a corpus.
The multilingual MS MARCO (mMARCO, Bonifacio et al., 2021) is a multilingual version of the well-known MSMARCO dataset (Bajaj et al., 2016), translated into Spanish using both open-source machine translation and Google Translate. Even if not originally in Spanish, this translated dataset is a substantial resource for training IR models since it leverages the large set of questions sampled from Bing’s search query logs which make up the original dataset. For each question, annotators judged the relevance of around 10 passages retrieved by Bing’s IR system. The collection of documents is the union of all retrieved passages.
A non-translated dataset is the Spanish Question Answering Corpus (SQAC, Gutiérrez-Fandiño et al., 2022). Inspired by SQUAD (Rajpurkar et al., 2016), Spanish-speaking annotators were shown passages from the Spanish Wikipedia, Wikinews, and the AnCora corpus, and were asked to come up with up to five questions for each fragment, encouraging the use of synonyms to prevent lexical overlap. Even if it is not a true IR dataset because it lacks a collection of documents, the set of all relevant passages can be used as a corpus for IR.
MIRACL (Multilingual Information Retrieval Across a Continuum of Languages, Zhang et al., 2023) is a multilingual dataset that includes Spanish and 17 other languages, focusing on good-quality relevance judgments by native speakers. Annotators were given a random paragraph from a Wikipedia article and were asked to formulate related questions, which may or may not be answerable by the passage. For each question, other annotators assessed the relevance of 10 candidate passages retrieved by three different IR models. Although MIRACL contains a relatively small number of queries, it benefits from having more annotations per query than other datasets, and from the inclusion of negative relevance judgments as well.
In addition to these multi-domain datasets, we identified two datasets focusing on specific domains. The Spanish Passage Retrieval dataset (PRES, Kamateri et al., 2019) contains documents representing articles from reliable Spanish health websites. The dataset is made up of health-related queries and employs a pooling method for relevance annotations, whereby human annotators assess the relevance of the pooled, top-ranked documents retrieved by different IR systems. Each document is a health-related article.
Finally, the Multilingual European Parliament Dataset (Multi-EuP, Yang et al., 2023) leverages official debate topics from the European Parliament. Each topic is a query, and the documents are speeches from Spanish parliamentarians that took part in the debate. The relevance judgments indicate whether a speech was part of the debate on the given topic or not.
3 MessIRve
MessIRve is a new large-scale dataset for Spanish IR, designed to better capture the information needs of Spanish speakers across different countries. In this section we describe the data collection process (section 3.1) and the dataset splits (section 3.2), and perform a quality assessment (section 3.3).
3.1 Data collection
MessIRve is built by using questions from Google’s autocomplete API (www.google.com/complete) as queries, and answers from Google Search “featured snippets” that link to Wikipedia as relevant documents. This data collection strategy is inspired by the GooAQ dataset (Khashabi et al., 2021).
Queries
Starting with 120 predefined prefixes such as “qué” (what), “cómo” (how), “dónde” (where), “quién” (who), etc., we used the Google autocomplete API to obtain popular queries in Spanish starting with these prefixes (refer to Appendix A, Table 3 for the list of prefixes). We then iteratively expanded the set of prefixes based on the returned results until we reached a predefined number of results. The API does not always return queries that begin strictly with the provided prefixes, resulting in a dataset where not all queries begin with the initial prefixes.
We ran this process for 20 countries with Spanish as an official language, plus the United States, by adjusting the API parameters to return results in Spanish for each country. Equatorial Guinea was the only country left out because it doesn’t have a Google domain. The extraction took place between March and April 2024, so the queries may reflect popular interests during that period.
During the process we noticed that some API results were independent of the country-specific domain, so we obtained many queries that were not specific of any country. These queries are included in the dataset under the country label none.
Relevant documents
To find relevant passages for each query, we obtained “featured snippets” from Google Search that sourced from Wikipedia. These snippets are text excerpts displayed at the top of the search results page in response to a query, deemed by Google to be particularly relevant to the user’s information need. In 2018 these snippets were reported to have a 97.4% accuracy rate in terms of providing relevant information. [^4] https://blog.google/products/search/reintroduction-googles-featured-snippets/
To obtain snippets from Wikipedia, we appended the term “wikipedia” to each query before querying Google Search. We then extracted the entire Wikipedia paragraph associated with the featured snippet as the relevant document for each query. Queries not returning featured snippets from Spanish Wikipedia were discarded.
Corpus
We built a Spanish Wikipedia corpus with each paragraph corresponding to a single document. For this we used WikiExtractor (Attardi, 2015) to process the Spanish Wikipedia dump of 2024-04-01 and matched the extracted Wikipedia passages to the valid (query, document) pairs obtained from Google Search featured snippets.
The choice of Wikipedia as the dataset’s collection was motivated by its accessibility, ease of processing, and the wide range of topics it covers.
3.2 Data splits
Train Test
Country Code #Q #Q #Unique Rel. Docs #Unique Rel. Articles Avg Q Length Avg Rel. Doc Length
- full 571 120 160 099 62 739 48 327 5.8 80.3
- none 358 014 99 385 43 490 34 745 5.8 81.1
Argentina ar 22 261 5 780 3 985 3 724 5.5 80.1
Bolivia bo 25 015 4 707 3 181 2 892 5.4 80.4
Chile cl 22 330 5 564 3 781 3 526 5.4 79.5
Colombia co 25 589 5 992 3 978 3 684 5.7 80.5
Costa Rica cr 23 605 5 747 4 111 3 839 5.5 80.5
Cuba cu 21 805 5 053 3 494 3 265 5.5 81.3
Dominican Rep. do 27 470 5 719 3 816 3 486 5.7 81.4
Ecuador ec 27 289 6 384 4 346 3 944 5.9 83.5
Spain es 23 115 7 509 5 205 4 913 5.6 80.1
Guatemala gt 23 311 4 290 2 959 2 726 5.6 80.5
Honduras hn 26 736 5 690 3 939 3 675 5.6 80.5
Mexico mx 31 957 8 400 5 618 5 198 5.6 79.7
Nicaragua ni 28 064 5 902 4 008 3 730 5.5 79.1
Panama pa 25 760 5 801 4 121 3 834 5.5 80.8
Peru pe 25 531 5 804 3 948 3 665 5.5 80.6
Puerto Rico pr 26 672 6 280 4 388 4 098 5.5 80.6
Paraguay py 24 717 5 474 3 867 3 574 5.5 78.6
El Salvador sv 25 920 5 821 4 078 3 785 5.6 79.6
United States us 23 197 4 535 2 942 2 784 5.6 81.1
Uruguay uy 20 856 5 571 4 062 3 816 5.4 77.4
Venezuela ve 26 730 6 126 4 091 3 715 5.7 82.8
Table 2: The MessIRve dataset. #: number of, Q: queries, Rel.: relevant, Doc.: document. Queries come from Google’s autocomplete API and documents are Wikipedia paragraphs. Therefore, a paragraph may be relevant to multiple queries. The full set includes queries that are not specific to any country (code=none) plus the set of unique queries from all countries. Each query has one relevant document in each country, while in the full set, a query may have multiple relevant documents. The statistics for the training set are nearly identical to those of the test set and are thus omitted.
The dataset is partitioned into training and test queries in such a way that the Wikipedia article to which the paragraph belongs is present in only one of the splits. The partitioning was done by country, with about 20% of the articles assigned to the test set. This split by article aims to reduce the overlap in topics between the training and test sets.
Queries not specific to any country (label none) were combined with the set of unique queries from all countries to form a full set, which is also split into training and test sets. Unlike the country-specific sets, in the full set some queries can have multiple relevant documents because the same query may return different featured snippets in different country domains. Statistical information about the dataset splits is provided in Table 2.
3.3 Quality assessment
We evaluated the quality of the dataset using a method inspired by Rybak (2023), sampling query-document pairs from the full test set and manually rating pairs on three binary criteria:
  1. 1.
    Query correctness: Can the information need of the user be understood? This criterion can be considered met even if the query is not grammatically correct, which is the case for many queries issued to search engines. For example, the query “porque se llama bogota” (Why is it called Bogotá?) can be considered correct even though it is missing question marks, accents, capitalization, etc.
  2. 2.
    Query unambiguity: Can a good answer to the query be given without providing further detail? Raters were instructed to consider that the query may only make sense in the context of the user’s country, which may not be explicitly stated. For example, the query “cómo se llama el presidente” (what is the name of the president) is ambiguous without specifying the country, but it is clear that the user is asking about the president of their country, and so it can be considered unambiguous.
  3. 3.
    Document relevance: Does the document help answer the query? Even if the passage does not provide an explicit and thorough answer, it can be considered relevant if it contains information that, when combined with other sources, could help answer the query.
Ratings were performed by four unpaid native Spanish speakers from Argentina. Each rater evaluated 200 query-document pairs, with 50 pairs being common across all raters. This resulted in a total of 650 unique rated pairs.
93.5% of the queries were rated as correct, 90.1% were rated as unambiguous, and 90.2% of the documents were considered relevant to their queries, using the majority vote for overlapping queries. These results show that the dataset’s quality is acceptable, as the queries are generally clear and the documents annotated as relevant are most likely to contain information that helps answer the query.
Because these three criteria are intrinsically ambiguous, i.e. they depend on each rater’s interpretation, we calculated the agreement amongst raters by computing the percentage of the 50 common queries on which all annotators agree. The percentages are: 78% for query correctness, 74% for query unambiguity, and 78% for document relevance. These results support the reliability of the quality assessment.
4 Comparison with existing datasets
A key feature of MessIRve is the detailed documentation of its data collection process. Existing open-domain datasets, such as MIRACL and SQAC, even if they are relatively well-documented, lack detailed descriptions of how their Wikipedia passages were selected before annotation.
MessIRve is also the only dataset that aims at explicitly accounting for the diverse varieties of Spanish spoken in different countries. This contrasts with all other datasets, which either do not consider dialectal variations or do not provide clear information about their inclusion.
To validate the dataset’s dialectal diversity, we listed the most distinctive words from each country’s queries according to a frequency-based log odds ratio score (see Table 4 in Appendix A). Each country’s queries include terms that are specifically relevant to that country, such as “green” for the United States (referring to the Green Card), “checar” for Mexico (meaning “to check”), “frailejones” for Colombia (a type of plant), “euskera” for Spain (the Basque language), and “CUIT” for Argentina (a tax identification number), when considering the five countries with the most speakers.
Moreover, MessIRve is substantially larger than other datasets (see Table 1). The mMARCO dataset, while large, involves translations from English, thus introducing potential translation errors and artifacts which may affect the quality of the dataset. Moreover, the queries in mMARCO are derived from English-speaking users’ queries, which may not be as relevant to Spanish-speaking users (see section 4.1 for a more detailed analysis).
One feature of our dataset is the sparsity of its judgments, with only about one relevance judgment per query, similar to the widely used MS MARCO dataset. This means that, as compared to datasets with more judgments per query, such as MIRACL, it is more likely that relevant passages are not identified as such, and that IR systems will retrieve more false negatives, that is, passages erroneously considered non-relevant.
Similarly, unlike MIRACL, our dataset does not include passages explicitly judged as negative (non-relevant). This is a limitation for the training of models that require negative examples, and implies that negative mining techniques would be needed to train models with MessIRve.
Nevertheless, the use of sparse, only-positive judgments is a common practice in IR research due to the challenges and high costs associated with obtaining more complete annotations. Previous research has shown that findings based on sparse judgments, like those of MS MARCO, align well with results from more comprehensive and costly judgments, suggesting that while more exhaustive annotations may change absolute scores, they are unlikely to affect qualitative conclusions regarding the ranking of IR systems (Zhang et al., 2021a).
The automated data collection approach was primarily driven by the need for a large-scale dataset that did not require the costly human annotations associated with datasets like MIRACL and MS MARCO. This approach sacrifices control over quality because it relies on both the reliability of the automatically extracted queries and the accuracy of Google Search featured snippets. While this method may be seen as less dependable than human annotations, we have conducted a quality assessment to ensure the dataset’s reliability, as described in section 3.3.
In the following section we perform a comparison between the datasets in terms of their topics.
4.1 Analysis of topics
MessIRve’s scale allows for a wider variety of topics than other datasets. For example, the full set of MessIRve contains judged passages coming from 84,284 unique Wikipedia articles, while MIRACL covers 17,640 articles and SQAC, 3,823 articles (considering training, validation and test sets).
Figure 1: Topics in Spanish IR datasets. Topics are sorted by overall frequency from top to bottom, and are labeled with an ID and the four most representative words according to their c-TF-IDF values (see Grootendorst, 2022). Numbers in the heatmap represent the % of queries in a given dataset that belong to a given topic. The color scale is shared across datasets. For the sake of clarity, the names of topics were translated from Spanish to English by the authors.
To more thoroughly analyze the topics covered by the queries of the different datasets, we employed BERTopic with PCA of 30 dimensions of multilingual-e5-large embeddings, and KMeans clustering with 30 clusters (Grootendorst, 2022; Wang et al., 2024). We assume just one topic per query, and focus only on the queries of the general-domain datasets: MIRACL, SQAC, mMARCO, and MessIRve full. For every dataset we use the complete set of available queries.
The clusters were labeled with the four most representative words according to their c-TF-IDF values (Grootendorst, 2022). To provide a summary of a particular topic in a given dataset, representative queries were found by sampling a subset of queries from the dataset belonging to the topic and calculating which ones were closest to the topic’s c-TF-IDF representations based on cosine similarity. Representative words of a topic in a dataset were also found by considering the highest c-TF-IDF scores of the words from the dataset.
The distribution of topics accross datasets (Figure 1) reveals two broad miscellaneous topics which are relatively prevalent in all datasets: queries concerning entertainment celebrities, including those in sports, music, movies, and other media (topic 0) and general definitions (topic 1).
MIRACL and SQAC show a higher frequency of topics related to war and politics (topic 19), religion and the Bible (topic 24), and inventions or discoveries (topic 25). This can be attributed to their dataset construction methodology, which starts from Wikipedia passages rather than user queries. Compared to other datasets, MessIRve shows a prominence of queries related to users experiencing health-related pains (topic 2), while mMARCO features topics related to household economy (topics 5 and 6) and computers and technology (topic 8).
It is interesting to note that topic 23, relatively prevalent in mMARCO but not in MessIRve, focuses on US demographics. This topic features representative documents such as “dallas nc se encuentra en qué condado” (Dallas, NC is located in what county), “en qué condado se encuentra grove city ohio” (In what county is Grove City, Ohio located), and “cual es el condado de washington tx” (Which is the county of Washington, TX).
This contrasts with the queries from topic 4, which are relevant in MessIRve, SQAC, and MIRACL, and focus on demographic issues. Representative queries from this topic in MessIRve include “en donde queda ciudad del este” (where is Ciudad del Este), “donde queda logroño” (where is Logroño), and “por donde queda montañita” (where is Montañita). These queries are likely to be more representative of the information needs of Spanish speakers than those in topic 23 of mMARCO, which probably come from English-speaking users.
Similarly, the queries in topic 0 of mMARCO and MessIRve highlight differences in popular entities within Spanish-speaking countries. For example, a representative mMARCO query is “en qué equipo juega michael sam” (what team does Michael Sam play for), referring to a US NFL player, while one from MessIRve is “en cuál equipo juega james rodriguez” (which team does James Rodriguez play for), referring to a Colombian soccer player. This behaviour is expected because MessIRve queries are obtained from Google’s autocomplete API by country domains, while mMARCO queries are derived from English-speaking users’ queries and translated into Spanish.
Finally, the distribution of topics in MessIRve and mMARCO is more uniform than in MIRACL and SQAC, which have a few topics that are much more frequent than others. This is likely a result of the larger size of MessIRve and mMARCO, allowing for a wider variety of topics.
However, MIRACL and SQAC contain well-formed questions, while MessIRve includes more colloquial and informal questions, which better represent the queries that users might issue to a search engine. This feature is also present in mMARCO, but affected by translation quality. See Table 5 in Appendix A for further details on representative words and documents for the most frequent topics.
We stress that the topic analysis of MessIRve’s queries is not intended to provide a faithful representation of the real distribution of popular topics for the Spanish-speaking community. Instead, it is just a descriptive analysis of the dataset, showcasing the kinds of topics included and their relative importance as compared to other datasets.
5 Experiments
We evaluate various IR models on MessIRve’s test sets in a zero-shot manner. In section 5.1 we describe the experimental setup, and in section 5.2 we present the results.
5.1 Experimental setup
We consider the following models for information retrieval in Spanish:
BM25
(Robertson et al., 2009). BM25 is a popular lexical IR model based on the token overlap between query and documents. We index the corpus and run retrieval using Pyserini with default parameters and Spanish analyzer (Lin et al., 2021). Despite its simplicity, BM25 is a strong baseline for IR (Thakur et al., 2021).
MIRACL-mdpr-es
(Zhang et al., 2023). This is a bi-encoder initalized from multilingual Dense Passage Retriever (mDPR, Zhang et al., 2021b), which is first pre-fine-tuned using the training set of MS MARCO and further fine-tuned on the MIRACL Spanish training set. We use the mdpr-tied-pft-msmarco-ft-miracl-es checkpoint. [^5] https://huggingface.co/castorini/mdpr-tied-pft-msmarco-ft-miracl-es As far as we know, this is the only well-documented, Spanish-specific retriever available.
E5-large
(Wang et al., 2024). This bi-encoder is initialized from the multilingual model XLM-R (Conneau et al., 2020) and then trained with weak supervision in a mixture of multilingual text pairs, followed by supervised fine-tuning in annotated data which is mostly in English but includes some multilingual data e.g. MIRACL. We use the multilingual-e5-large checkpoint. [^6] https://huggingface.co/intfloat/multilingual-e5-large
OpenAI-large
. Although the architecture and training data of this model are not publicly disclosed, previous models (Neelakantan et al., 2022) consist of a single encoder that maps documents and query to embeddings, and is trained with contrastive learning across several supervised datasets. The model we use, text-embedding-3-large, is accessible via an API. [^7] https://platform.openai.com/docs/guides/embeddings
In all cases, we append the title of the Wikipedia article to the document text before retrieval to provide additional context.
We use two standard metrics to measure retrieval performance:
  • Recall@100: the fraction of relevant documents within the top 100 results, averaged over all queries.
  • nDCG@10: the normalized Discounted Cumulative Gain. It compares the rank of the top 10 results to the ideal ranking where relevant documents are ranked higher. It is averaged over all queries.
All evaluations are conducted in a zero-shot setting, where the models are not fine-tuned on any training set. We leave this for future work.
5.2 Results
Figure 2: Baseline results on MessIRve test sets.
The OpenAI-large and E5-large models outperform smaller models (MIRACL-mdpr-es) and the lexical model (BM25) across all dataset partitions (see Figure 2). In terms of nDCG@10, OpenAI-large achieves an average score in the full set of .433 compared to E5-large’s .430, indicating similar ranking effectiveness. However, for Recall@100, OpenAI-large consistently outperforms E5-large, with a recall of .899 versus .847.
BM25 consistently performs the worst across all subsets (nDCG@10=.167 and Recall@100=.530 in the full set), highlighting the limitations of lexical retrieval methods compared to dense models.
Although MIRACL-mdpr-es is specifically fine-tuned for Spanish, it still lags behind the larger models, with nDCG@10 and Recall@100 scores of .259 and .626, respectively, suggesting that model size and fine-tuning strategy might play a more critical role than language-specific tuning. For example, while MIRACL-mdpr-es has 178M parameters and is pre-fine-tuned with MS MARCO and then fine-tuned with Spanish MIRACL, E5-large has 560M parameters, is pre-trained in a mixture of multilingual text pairs, and then fine-tuned in at least 10 supervised datasets, including MS MARCO and multilingual data like MIRACL itself.
Finally, the three dense models follow a similar performance pattern, with the highest scores in the bo (Bolivia) and gt (Guatemala) subsets and the lowest in ar (Argentina) and mx (Mexico).
6 Discussion and conclusions
We presented MessIRve, a large-scale dataset for information retrieval in Spanish. MessIRve queries originate from Spanish speakers from almost all Spanish-speaking countries, ensuring relevance to the information needs of the Spanish-speaking community. This contrasts with other datasets, which are either translated from English or do not consider dialectal variations. Moreover, the dataset’s large scale allows it to cover a wide variety of topics, unlike smaller datasets.
We performed a quality assessment to ensure the reliability of the dataset, and we provided a detailed comparison with existing Spanish IR datasets, highlighting the dataset’s strengths and limitations. We also provide baseline results of the performance of various IR models on the dataset.
It must be noted that the dataset was collected using an IR system in itself, Google’s “featured snippets”. Although there’s no certainty about how Google’s IR system works, it is likely that it considers many other factors besides the text of the query and the documents, such as inbound links, user behavior, the structure of the web, etc., implying significant computational resources. This probably contributes to the high quality of the dataset, and so makes it a valuable resource for training and evaluating open-source IR models. However, it also introduces an evaluation bias towards Google’s own IR system, as the dataset may not be used to identify IR systems that are better than Google’s own system.
In this work, we have focused on zero-shot baselines. Future work includes fine-tuning the models on the training set of MessIRve, evaluating models on other Spanish IR datasets, and creating a unified Spanish IR benchmark with all datasets.
Limitations
The dataset collected may not be representative of the distribution of popular topics in each country, as it only includes a sample of popular queries in Spanish-speaking countries at the time of collection. Moreover, it only consists of queries which have an answer in Spanish Wikipedia, and that Google is confident in answering.
The quality assessment of the dataset was performed by four of the authors, which may introduce bias. This was done due to the lack of resources to hire external raters.
The dataset should not be used for commercial purposes.
Acknowledgments
This work used computational resources from CCAD – Universidad Nacional de Córdoba (https://ccad.unc.edu.ar/), which are part of SNCAD – MinCyT, República Argentina, and from NodaIA San Francisco (Ministry of Science and Technology of the Province of Córdoba, Argentina).
We thank OpenAI for their financial support to run experiments with their models. We also thank authors of GooAQ (Khashabi et al., 2021) and Pyserini (Lin et al., 2021) for publicly sharing their code, which significantly contributed to our research.
Appendix A Supporting information
Table 3 shows the seed prefixes used to obtain the queries for each country in the MessIRve dataset. These are loosely based on Khashabi et al. (2021). We include incorrect, non-accented, versions of words, e.g. “como” instead of “cómo”, to account for common spelling mistakes in Spanish queries.
Table 4 shows the most distinctive words for each country in the MessIRve dataset. These words are obtained by calculating the log odds ratio of the frequency of each word in the queries for each country with respect to the queries of the full subset, and selecting the words with the highest scores. The words are indicative of the dialectal variety of Spanish spoken in each country.
Table 5 shows the most representative words and queries for the 8 most frequent topics in each dataset. Representative words and queries were selected based on the c-TF-IDF scores (Grootendorst, 2022).
‘‘a cual’’, ‘‘a cuales’’, ‘‘a cuantas’’, ‘‘a cuantos’’, ‘‘a cuál’’, ‘‘a cuáles’’, ‘‘a cuántas’’, ‘‘a cuántos’’, ‘‘a donde’’, ‘‘a dónde’’, ‘‘a que’’, ‘‘a quien’’, ‘‘a quienes’’, ‘‘a quién’’, ‘‘a quiénes’’, ‘‘a qué’’, ‘‘adonde debe’’, ‘‘adónde debe’’, ‘‘ante cual’’, ‘‘ante cuales’’, ‘‘ante cuál’’, ‘‘ante cuáles’’, ‘‘ante que’’, ‘‘ante quien’’, ‘‘ante quienes’’, ‘‘ante quién’’, ‘‘ante quiénes’’, ‘‘ante qué’’, ‘‘bajo cual’’, ‘‘bajo cuales’’, ‘‘bajo cuál’’, ‘‘bajo cuáles’’, ‘‘bajo que’’, ‘‘bajo qué’’, ‘‘buenas razones para’’, ‘‘buenos motivos para’’, ‘‘como’’, ‘‘con cual’’, ‘‘con cuales’’, ‘‘con cuanta’’, ‘‘con cuantas’’, ‘‘con cuanto’’, ‘‘con cuantos’’, ‘‘con cuál’’, ‘‘con cuáles’’, ‘‘con cuánta’’, ‘‘con cuántas’’, ‘‘con cuánto’’, ‘‘con cuántos’’, ‘‘con que’’, ‘‘con quien’’, ‘‘con quienes’’, ‘‘con quién’’, ‘‘con quiénes’’, ‘‘con qué’’, ‘‘cual’’, ‘‘cuales’’, ‘‘cuando’’, ‘‘cuanta’’, ‘‘cuantas’’, ‘‘cuanto’’, ‘‘cuantos’’, ‘‘cuál’’, ‘‘cuáles’’, ‘‘cuándo’’, ‘‘cuánta’’, ‘‘cuántas’’, ‘‘cuánto’’, ‘‘cuántos’’, ‘‘cómo’’, ‘‘de cual’’, ‘‘de cuanta’’, ‘‘de cuantas’’, ‘‘de cuanto’’, ‘‘de cuantos’’, ‘‘de cuál’’, ‘‘de cuánta’’, ‘‘de cuántas’’, ‘‘de cuánto’’, ‘‘de cuántos’’, ‘‘de donde’’, ‘‘de dónde’’, ‘‘de que’’, ‘‘de quien’’, ‘‘de quienes’’, ‘‘de quién’’, ‘‘de quiénes’’, ‘‘de qué’’, ‘‘deberia’’, ‘‘deberian’’, ‘‘desde donde’’, ‘‘desde dónde’’, ‘‘desde que’’, ‘‘desde qué’’, ‘‘donde’’, ‘‘durante cual’’, ‘‘durante cuales’’, ‘‘durante cuanto’’, ‘‘durante cuantos’’, ‘‘durante cuál’’, ‘‘durante cuáles’’, ‘‘durante cuánto’’, ‘‘durante cuántos’’, ‘‘durante que’’, ‘‘durante qué’’, ‘‘dónde’’, ‘‘en cual’’, ‘‘en cuales’’, ‘‘en cuantas’’, ‘‘en cuanto’’, ‘‘en cuantos’’, ‘‘en cuál’’, ‘‘en cuáles’’, ‘‘en cuántas’’, ‘‘en cuánto’’, ‘‘en cuántos’’, ‘‘en donde’’, ‘‘en dónde’’, ‘‘en que’’, ‘‘en quien’’, ‘‘en quién’’, ‘‘en qué’’, ‘‘entre cuales’’, ‘‘entre cuáles’’, ‘‘entre que’’, ‘‘entre quienes’’, ‘‘entre quiénes’’, ‘‘entre qué’’, ‘‘hacia cual’’, ‘‘hacia cuál’’, ‘‘hacia que’’, ‘‘hacia quien’’, ‘‘hacia quién’’, ‘‘hacia qué’’, ‘‘hasta cual’’, ‘‘hasta cuando’’, ‘‘hasta cuanta’’, ‘‘hasta cuantas’’, ‘‘hasta cuanto’’, ‘‘hasta cuantos’’, ‘‘hasta cuál’’, ‘‘hasta cuándo’’, ‘‘hasta cuánta’’, ‘‘hasta cuántas’’, ‘‘hasta cuánto’’, ‘‘hasta cuántos’’, ‘‘hasta donde’’, ‘‘hasta dónde’’, ‘‘hasta que’’, ‘‘hasta qué’’, ‘‘mediante cual’’, ‘‘mediante cuales’’, ‘‘mediante cuál’’, ‘‘mediante cuáles’’, ‘‘mediante que’’, ‘‘mediante qué’’, ‘‘motivos de por que’’, ‘‘motivos de por qué’’, ‘‘motivos de’’, ‘‘motivos del’’, ‘‘motivos para’’, ‘‘motivos por los que’’, ‘‘motivos por los qué’’, ‘‘para cual’’, ‘‘para cuantas’’, ‘‘para cuantos’’, ‘‘para cuál’’, ‘‘para cuántas’’, ‘‘para cuántos’’, ‘‘para donde’’, ‘‘para dónde’’, ‘‘para que’’, ‘‘para quien’’, ‘‘para quienes’’, ‘‘para quién’’, ‘‘para quiénes’’, ‘‘para qué’’, ‘‘podria’’, ‘‘por cuanta’’, ‘‘por cuantas’’, ‘‘por cuanto’’, ‘‘por cuantos’’, ‘‘por cuánta’’, ‘‘por cuántas’’, ‘‘por cuánto’’, ‘‘por cuántos’’, ‘‘por donde’’, ‘‘por dónde’’, ‘‘por que debe’’, ‘‘por que deberia’’, ‘‘por que no’’, ‘‘por que’’, ‘‘por qué debe’’, ‘‘por qué deberia’’, ‘‘por qué no’’, ‘‘por qué’’, ‘‘porque’’, ‘‘porqué’’, ‘‘pros y contras de’’, ‘‘puede’’, ‘‘pueden’’, ‘‘que’’, ‘‘quien’’, ‘‘quienes’’, ‘‘quién’’, ‘‘quiénes’’, ‘‘qué’’, ‘‘razones de por que’’, ‘‘razones de por qué’’, ‘‘razones de’’, ‘‘razones del’’, ‘‘razones para’’, ‘‘razones por las que’’, ‘‘razones por las qué’’, ‘‘segun quien’’, ‘‘segun quienes’’, ‘‘segun quién’’, ‘‘segun quiénes’’, ‘‘según quien’’, ‘‘según quienes’’, ‘‘sobre cual’’, ‘‘sobre cuál’’, ‘‘sobre que’’, ‘‘sobre quien’’, ‘‘sobre quienes’’, ‘‘sobre quién’’, ‘‘sobre quiénes’’, ‘‘sobre qué’’, ‘‘tras cuantas’’, ‘‘tras cuanto’’, ‘‘tras cuantos’’, ‘‘tras cuántas’’, ‘‘tras cuánto’’, ‘‘tras cuántos’’
Table 3: Seed prefixes used to obtain queries from the Google autocomplete API.
ar antula 1982 cuit pampita renga rivadavia
bo samaipata bulo pagador willka katari evo
cl huilo pichilemu pasamos apuros chiloe colocolo
co eliecer gaitan covenas nevados frailejones caldas
cr puntarenas alajuela bifonazol guanacaste rica hidroxizina
cu maceo cienfuegos baragua monilia yate bayamo
do enriquillo bani balaguer luperon dominicano altagracia
ec montubios veintimilla guayasamin proponen ecuatoriana roldos
es croquetas declinaciones protestan supervivientes naiara euskera
gt mazatenango monterrico pacaya tikal verapaz xincas
hn sula tegucigalpa morazan copan lempira yojoa
mx maluma decanato redactaron kilataje kumbia checar
ni masaya gueguense matagalpa sandino managua esteli
pa delfia panamena cortez rolo martinelli torrijos
pe boluarte quispe chiclayo benavides chimbote pisco
pr decadron scan micoplasma coqui baclofen mycoplasma
py paraguaya itaipu solano plaqueta paraguay paraguayo
sv pital parlacen siguanaba lempa fulcro salvador
us taxes apocalypto millions contenedor million green
uy artigas penarol uruguaya montevideo sambayon sapitos
ve camejo hipolita morrocoy avanzadora ribas cardenales
Table 4: Words in the queries of MessIRve are representative of the dialectal varieties of Spanish and the topics of interest in each country. To find distinctive words for each country, we calculated the log odds ratio of the frequency of each word in each country with respect to the queries of the full subset, using smoothing with a prior of 0.1 (Jurafsky and Martin, 2024). We consider only words with at least 10 occurrences in the country subset and show the top 6 words with the highest log odds ratio for each country.
messIRve mMARCO MIRACL SQAC
Topic 2 Topic 5 Topic 19 Topic 0
n=61,797 (8.8%) n=60,499 (7.4%) n=498 (11.5%) n=4,314 (22.9%)
duele, razones, motivos, da, salen costo, cuesta, salario, promedio, precio paises, guerra, nazi, pais, imperio cargo, profesion, equipo, partido, dirige
porque da dolor y ardor en la boca del estomago costo promedio de la universidad ¿Cuándo ocurrió la primera guerra mundial? ¿En qué equipo juega Ronaldo?
Topic 3 Topic 1 Topic 0 Topic 19
n=48,188 (6.9%) n=59,825 (7.3%) n=455 (10.5%) n=2,693 (14.3%)
consiste, derechos, importante, sociales, cuales significa, empresa, hace, archivo, red cantante, banda, llama, personaje, pelicula guerra, gobierno, pais, presidente, mundial
porque son importante los derechos humanos la empresa de redes más rica del mundo ¿Qué actor interpretó a Ray Charles en el cine? ¿En qué continente tuvo lugar la Guerra de la Independencia Argentina?
Topic 4 Topic 6 Topic 25 Topic 4
n=45,908 (6.5%) n=47,105 (5.8%) n=380 (8.8%) n=1,733 (9.2%)
queda, region, pertenece, continente, rio impuestos, credito, obtener, licencia, seguro primer, primera, invento, fundo, creo lugar, adonde, halla, encuentra, ciudad
en qué continente queda australia ¿Cuánto tiempo se tarda en obtener un reembolso de impuestos? ¿Dónde se inventó el baloncesto? ¿Con qué territorio limita Dubái por su parte occidental?
Topic 0 Topic 0 Topic 24 Topic 24
n=43,578 (6.2%) n=46,039 (5.6%) n=319 (7.4%) n=1,485 (7.9%)
murio, juega, canta, gano, equipo murio, canta, elenco, interpreta, pelicula biblia, historia, mitologia, religion, dios obra, obras, segun, jesus, iglesia
con quien canta michael jackson todo mi amor eres tu quien canta la cancion uno ¿Quién era David en la Biblia? ¿Quiénes escribieron los evangelios que se creen bastante cercanos a la época en que vivió Jesús de Nazaret?
Topic 11 Topic 9 Topic 4 Topic 11
n=43,539 (6.2%) n=44,004 (5.4%) n=313 (7.2%) n=1,139 (6.1%)
cuantas, cuantos, existen, municipios, jugadores sintomas, enfermedad, causar, efectos, enfermedades paises, capital, pais, ciudad, ciudades personas, cuantas, cuantos, victimas, fallecieron
cuántos idiomas indígenas existen en méxico ¿Qué enfermedad causa dolor de estómago, diarrea y es contagiosa? ¿En qué región se encuentra la República de las Islas Marshall? ¿Cuántas personas utilizan este servicio en todo mundo?
Topic 7 Topic 8 Topic 13 Topic 20
n=36,811 (5.2%) n=43,946 (5.4%) n=312 (7.2%) n=978 (5.2%)
sirve, tomar, medicamento, mg, cura numero, telefono, windows, computadora, excel botanica, cuales, vegetacion, populares, caracteristicas produjo, comicios, competicion, lugar, partido
que pastilla puedo tomar para el dolor de cabeza Excel número de teléfono de servicio al cliente ¿Cuántos colores puede reconocer un perro? ¿Cómo se pasa del horario de invierno al de verano?
Topic 12 Topic 14 Topic 3 Topic 25
n=30,631 (4.4%) n=43,221 (5.3%) n=261 (6.0%) n=961 (5.1%)
energia, agua, tierra, ambiente, planeta definicion, definir, significa, medica, biologia principales, estudia, diferencia, cuales, paleontologos primer, primera, escribio, creo, descubrio
qué tipo de movimiento realiza el planeta tierra alrededor del sol definición de sobre ¿Qué habilidades debe tener un director ejecutivo? ¿Quién descubrió los rayos X?
Topic 1 Topic 23 Topic 15 Topic 1
n=28,177 (4.0%) n=38,566 (4.7%) n=240 (5.5%) n=636 (3.4%)
trata, consiste, marca, juego, gama condado, encuentra, poblacion, tx, ca botanica, significan, significa, idioma, sigla siglas, dedica, drcafta, posicion, objetivo
de que trata te para 3 ¿En qué condado se encuentra la ciudad de texas, tx? ¿Qué quiere decir la palabra pseudo? ¿Cómo se llama el programa?
Table 5: Representative words and queries from topics in Spanish IR datasets. These were selected based on the c-TF-IDF scores of the words and queries in each topic, respectively. Refer to section 4.1 for more details.